Q.FIND THE EQUATION OF THE CIRCLE TOUCHING THE BOTH THE COORDINATE AXES AND PASSING THROUGH THE POINT (3,-6)?

Grade:12th pass

3 Answers

vikas askiitian expert
13 years ago

general equation of circle is

(x-h)2 + (y-k)2 = R2

when circle touches both the axis then h = k = R

(x-h)2 + (y-h)2 = h2

now this circle passes through (3,-6) so

(3-h)2 + (-6-h)2 = h2

 h2 +6h+45 = 0

 descriminant = -ve , hence h is imaginary so there will be no circle under this condition....

Fawz Naim
13 years ago

Let the radius of the circle be r. Then as the circle touches both the coordinate axes, the coordinates of its centre are (r,r)

and its distance from the point (3,-6) is the radius of the circle

(r-3)^2+(r+6)^2=r^2

now calculate it

Paras Arora
7 years ago
the answer us x^2+y^2-30x+30y+225.
 
this can be done by assuming pt be (h,h). and solving
 
circle equation.